find the total differential of the function \\(f(x,y)=ye^{x+y}\\)
1
Expert's answer
2019-10-21T11:01:15-0400

"f(x,y)=ye^{(x+y)}"

Total differential of the function is "d f =(\frac{\partial f}{\partial x})dx+(\frac{\partial f}{\partial y})dy"

"d f=\frac{\partial }{\partial x}(ye^{(x+y)})dx+\frac{\partial }{\partial y}(ye^{(x+y)})dy"

"d f=ye^{(x+y)}dx" "+(ye^{(x+y)}+e^{(x+y)})dy"